Find the two consecutive integers that have a sum of 39.


1
Expert's answer
2021-09-28T00:02:57-0400

Let the two consecutive integers be n and (n+1)

"n+(n+1)=39"

"2n=38"

"n=19"

"n+1=19+1=20"

So the two integers are 19 and 20